OpenStudy (anonymous):

STUCK FOR DAYS!!!! NEED HELP ASAP!!!!! MEDALS!!!!! FAN-ING!!!!! WHATEVER!!! I'LL DO ANYTHING FOR HELP!!!!!!!

OpenStudy (anonymous):

Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days Stock A B C Price f(x) = 25(1.08)x x F(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Step 1: Write equations for each Stock (fill in the blanks) Stock A B C D Values for D Starting price 25 22 30 Second lowest starting price: 23 Rate of growth 8% each day 4% each day Second highest rate of growth: 6% Equation f(x) = 25(1.08)x F(x)=26(1.3)^x F(x)=30(1.4)^x   Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days Stock A B C Price f(x) = 25(1.08)x x F(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Step 1: Write equations for each Stock (fill in the blanks) Stock A B C D Values for D Starting price 25 22 30 Second lowest starting price: 23 Rate of growth 8% each day 4% each day Second highest rate of growth: 6% Equation f(x) = 25(1.08)x F(x)=26(1.3)^x F(x)=30(1.4)^x  

OpenStudy (anonymous):

what the question

OpenStudy (anonymous):

ik it seems like alot but i really need help

OpenStudy (anonymous):

sorry man I don't know how to do this

OpenStudy (anonymous):

its ok do u know any one that can help me ?

OpenStudy (triciaal):

stop and breathe

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

done XD

OpenStudy (anonymous):

Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days Stock A B C Price f(x) = 25(1.08)x x F(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Step 1: Write equations for each Stock (fill in the blanks) Stock A B C D Values for D Starting price 25 22 30 Second lowest starting price: 23 Rate of growth 8% each day 4% each day Second highest rate of growth: 6% Equation f(x) = 25(1.08)x F(x)=26(1.3)^x F(x)=30(1.4)^x  

OpenStudy (anonymous):

mb i pasted it 2 times

OpenStudy (triciaal):

really don't even want to read something this long

OpenStudy (triciaal):